OpenStudy (anonymous):

integral of sin^2(x)cos(3x)dx

OpenStudy (aum):

Find cos(3x) in terms of cos(x). cos(3x) = cos(2x + x) = ?

OpenStudy (anonymous):

Thanks, its cos(2x)cos(x)-sin(2x)sin(x), right? If so I think i got it from here

OpenStudy (aum):

Correct. Go further and expand cos(2x) and sin(2x).

OpenStudy (anonymous):

I got that to become 4cos^3(x)-3cos(x) which then results in int sin^2(x)4cos^3(x) -int sin^2(x)cos(x)

OpenStudy (aum):

sin^2(x) * { 4cos^3(x) - cos(x) } = sin^2(x) * { 4cos^2(x) - 1 } * cos(x) = sin^2(x) * { 4(1-sin^2x) - 1 } * cos(x) = sin^2(x) * { 3 - 4sin^2x } * cos(x) = Let u = sin(x) u^2 * (3 - 4u^2) * du

OpenStudy (aum):

I have not checked if cos(3x) = 4cos^3(x) - cos(x) that you got is correct.

OpenStudy (anonymous):

I believe it is, i just finished the problem and everything worked out so hopefully that was correct Thanks a lot

OpenStudy (aum):

You are welcome.